Maria plans to take a trip in three years. She knows the total cost of the trip is $5,500. She has two investment options that will help her pay for the trip.   
a. Investment A: Earns 6.5%
b. Investment B: Earns 8%
i. How much would she have to invest in either investment to have the money she needs for her trip?
ii. Show your work.
Answer #1

We know that the present value=future value/(1+rate)^time

Hence,
Amount for A=5500/1.065^3=4553.170005
Amount for B=5500/1.08^3=4366.077326